OpenStudy (anonymous):

Twice the sum of a number and 7 equals 6

OpenStudy (potatoes.ramu):

Let the number be x

OpenStudy (potatoes.ramu):

So the equation you get is 2(x+7)=6

OpenStudy (potatoes.ramu):

And then you can solve for x! :)

OpenStudy (ybarrap):

$$ \text{Twice}(\color{red}{2\times})\text{ the sum}(\color{red}{+})\text{ of a number (call it }\color{red}{x}\text{) and 7 equals(}\color{red}{=}\text{) 6.}\\\\ \text{In symbols:}\\ 2\times(x+7)=6\\ $$

OpenStudy (anonymous):

yes ok